# SMART: Submodular Data Mixture Strategy for Instruction Tuning

### v1

@@ 1 Introduction
  • (RQ1) Given a huge multitask instruction-tuning dataset and a limited fine-tuning budget which is defined by the total number of $(prompt,response)$ instances that can be used for fine-tuning, how do we divide this budget among thousands of tasks present in the dataset? i.e., how many instances to sample from each task? and which instances to sample from each task?
@@
As a first step towards addressing both of the above RQs, we first define a common subset selection problem (more formally stated in Section 3) as follows - Given a huge collection of $M$ instruction-formatted task datasets, a task budget $M^{\prime}\leq M$ and a total budget ($N^{\prime}$) of $(prompt,response)$ pairs, which $M^{\prime}$ tasks to select, how many instances to select from each of these $M^{\prime}$ tasks and which instances to select? Note that RQ1 is an instance of this problem where $M^{\prime}=M$.
@@ 4.2 Finetuning Procedure
We evaluate SMART on three 7B parameter LLMs: Llama-2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Mistral Jiang et al. (2023). We fine-tune each model for 1 epoch on a given data mixture with a learning rate of $52e-$ 5 for Llama-2 and Falcon, and $65e-$ 6 for Mistral. A batch size of $64$, weight decay of $0.01$, and cosine learning rate decay with a linear warmup for the initial $1\%$ training steps are employed. The experiments run on 8 NVIDIA A100-SXM4-80GB GPUs, utilizing Flash Attention Dao (2023) for memory efficiency and speeding up finetuning. Our code is open-sourced here
@@ 4.3 Baselines
We compare SMART with two baseline mixture strategies: Examples Proportional Mixture and Equal Mixture  (Raffel et al., 2020)
@@
$25000$ EPM (Baseline-1) 30.82 46 46.71 43.05 40.63 40.59 25.5 31.67 38.05
EM (Baseline-2) 30.33 44.99 43.81 40.55 39.02 40.08 20.29 29.46 36.27
@@
SMART (Ours) 36.51 52.85 54.49 50.79 47.53 46.73 20.1 31.75 42.99
$100000$ EPM (Baseline-1) 32.66 50.45 51.36 47.18 44.22 43.33 26.36 33.46 41.13
EM (Baseline-2) 36.03 50.69 52.53 47.1 45.57 44.4 25.18 33.55 42.11
SMART (Ours) 37.36 55.38 55.46 52.95 49.11 47.17 24.22 34.62 44.94
$200000$ EPM (Baseline-1) 34.88 53.85 54.31 50.18 47.06 43.56 26.56 33.81 43.25
EM (Baseline-2) 38.6 53.58 54.72 51.47 48.57 40.51 24.75 31.93 43.79
SMART (Ours) 39.2 57.29 58.71 55.01 51.32 44.87 24.33 33.92 46.32
$400000$ EPM (Baseline-1) 38.16 56.46 56.99 52.56 49.84 47.74 26.02 36.08 45.88
EM (Baseline-2) 39.43 55.85 57.59 53.65 50.49 46.61 24.49 34.89 46
SMART (Ours) 39.77 57.18 60.17 54.79 51.72 46.43 26.35 36.27 47.28
17,591,640 Full FLAN 2022 42.44 59.1 61.82 55.1 53.43 47.16 27.6 36.64 48.6
@@ 4.5 Addressing RQ1 ($M^{\prime}=M$)
RQ1 is an instance of the subset selection problem from Section  3, where $M^{\prime}=M$, allowing all tasks for finetuning; however we still have a constraint ($N^{\prime}$) on total number of $(prompt,response)$ pairs. We use Graph-Cut and Facility Location functions from Table 1 in Stage-1 (Section 3.1.1) and Stage-2 (Section 3.1.2) of SMART respectively. This choice of $f_{1}$ and $f_{2}$ is determined via grid search discussed in Section 4.7. Table 2 compares SMART with baselines (on evaluation benchmarks described in Section 4.4) by finetuning Llama-2-7b, on data mixtures generated by varying $N^{\prime}$ in $\{25000,50000,100000,200000,400000\}$. As can be inferred from the table, SMART consistently performs better than traditional mixtures and also performs close to finetuning on full collection.
@@ 4.8 Ablation Study: LLM
$N^{\prime}$ Data Mix. Mistral $*{}^{*}$
$25000$ EPM(Baseline-1) 52.69 23.08
EM(Baseline-2) 51.65 21.1
SMART 53.33 22.71
$50000$ EPM(Baseline-1) 53.75 19.85
EM(Baseline-2) 53.16 22.95
SMART 53.8 26.31
$100000$ EPM(Baseline-1) 54.21 20.86
EM(Baseline-2) 54.55 24.84
SMART 53.72 27.22
$200000$ EPM(Baseline-1) 55.08 24.69
EM(Baseline-2) 54.7 24.38
SMART 55.17 30.95
Table 4: Comparison of SMART with baselines for Mistral and Falcon.
@@ Appendix A Code and Data
Our code for creating data mixtures is open-sourced at https://github.com/kowndinya-renduchintala/SMART. The code development utilized open-source tools, primarily relying on the HuggingFace library for model training with PyTorch as the underlying framework. Both PyTorch and HuggingFace are licensed under permissive licenses, with PyTorch under the BSD license and HuggingFace under the Apache 2.0 license. Additionally, submodular optimization was performed using SUBMODLIB, which is an openly accessible library on GitHub at https://github.com/decile-team/submodlib under the MIT license.

### v2

@@ 1 Introduction
  • (RQ1) Given a huge multitask instruction-tuning dataset and a limited fine-tuning budget which is defined by the total number of $(prompt,response)$ pairs that can be used for fine-tuning, how do we divide this budget among thousands of tasks present in the dataset? i.e., how many instances to sample from each task? and which instances to sample from each task?
@@
As a first step towards addressing both of the above RQs, we first define a common subset selection problem (more formally stated in Section 3) as follows - Given a huge collection of $M$ instruction-formatted task datasets, a task budget $M^{\prime}\leq M$ and a total budget $N^{\prime}$, of $(prompt,response)$ pairs, which $M^{\prime}$ tasks to select, how many instances to select from each of these $M^{\prime}$ tasks and which instances to select? Note that RQ1 is an instance of this problem where $M^{\prime}=M$.
@@ 4.2 Finetuning Procedure
We evaluate SMART on three 7B parameter LLMs: Llama-2 Touvron et al. (2023), Falcon Almazrouei et al. (2023), and Mistral Jiang et al. (2023). We fine-tune each model for 1 epoch on a given data mixture with a learning rate of $52e-$ 5 for Llama-2 and Falcon, and $65e-$ 6 for Mistral. A batch size of $64$, weight decay of $0.1$, and cosine learning rate decay with a linear warmup for the initial $1\%$ training steps are employed. The experiments run on 8 NVIDIA A100-SXM4-80GB GPUs, utilizing Flash Attention Dao (2023) for memory efficiency and speeding up finetuning. Our code is open-sourced here.
@@ 4.3 Baselines
We compare SMART with three baseline mixture strategies: Examples Proportional Mixture, Equal Mixture  (Raffel et al., 2020) and FLANv2 Mixture (Longpre et al., 2023)
@@
FLANv2 Mixture (Baseline-3)
Following Longpre et al. (2023), 40%, 32%, 20%, 5% and 3% are assigned to the 5 submixtures of FLAN 2022 - FLANv1, T0, NIV2, CoT and Dialog respectively and instances are sampled randomly from each submixture.
@@
$25000$ EPM (Baseline-1) 30.82 46 46.71 43.05 40.63 40.59 25.55 31.67 38.05
EM (Baseline-2) 30.33 45.01 43.81 40.55 39.03 40.08 20.29 29.46 36.27
FLANv2 (Baseline-3) 30.55 49.08 50.14 45.09 42.47 43.79 25.73 33.33 39.84
@@
FLANv2 (Baseline-3) 31.85 52.54 51.71 47.54 44.6 43.13 26.13 33.01 41.27
SMART (Ours) 36.51 53.06 54.49 50.79 47.58 46.73 20.1 31.75 43.03
$100000$ EPM (Baseline-1) 32.66 50.6 51.37 47.18 44.25 43.38 26.36 33.48 41.16
EM (Baseline-2) 36.03 50.7 52.53 47.1 45.57 44.4 25.18 33.55 42.11
FLANv2 (Baseline-3) 36.1 54.1 55.6 50.44 47.82 44.44 26.22 33.42 43.68
SMART (Ours) 37.36 55.38 55.47 52.95 49.11 47.26 24.22 34.66 44.96
$200000$ EPM (Baseline-1) 35.19 54.64 54.75 50.58 47.53 45.25 26.57 34.52 43.79
EM (Baseline-2) 38.6 54.05 54.72 51.47 48.68 41.36 24.75 32.28 43.96
FLANv2 (Baseline-3) 34.76 55.02 56.65 50.13 47.76 44.39 27.68 34.73 44.02
SMART (Ours) 39.2 57.29 58.71 55.01 51.32 47.99 24.47 35.27 46.7
$400000$ EPM (Baseline-1) 38.16 56.53 56.99 52.56 49.85 48.72 26.04 36.49 46.01
EM (Baseline-2) 39.43 55.97 57.59 53.65 50.52 47.37 26.08 35.8 46.29
FLANv2 (Baseline-3) 37.46 56.79 58.42 51.4 49.71 49.11 26.56 36.97 46.05
SMART (Ours) 39.77 57.39 60.17 54.79 51.77 49.25 26.35 37.43 47.65
17,591,640 Full FLAN 2022 42.44 59.1 61.82 55.1 53.43 50.7 27.6 38.11 49.03
@@ 4.5 Addressing RQ1 ($M^{\prime}=M$)
RQ1 is an instance of the subset selection problem from Section  3, where $M^{\prime}=M$, allowing all tasks for finetuning; however we still have a constraint ($N^{\prime}$) on total number of $(prompt,response)$ pairs. We use Graph-Cut and Facility Location functions from Table 1 in Stage-1 (Section 3.1.1) and Stage-2 (Section 3.1.2) of SMART respectively. This choice of $f_{1}$ and $f_{2}$ is determined via grid search discussed in Section 4.7. Table 2 compares SMART with baselines (on evaluation benchmarks described in Section 4.4) by finetuning Llama-2-7b, on data mixtures generated by varying $N^{\prime}$ in $\{25000,50000,100000,200000,400000\}$. As can be inferred from the table, SMART consistently performs better than all the three baselines and also performs close to finetuning on full collection.
@@ Appendix A Code and Data
Our code for creating data mixtures is open-sourced at https://anonymous.4open.science/r/SMART-D829. The code development utilized open-source tools, primarily relying on the HuggingFace library for model training with PyTorch as the underlying framework. Both PyTorch and HuggingFace are licensed under permissive licenses, with PyTorch under the BSD license and HuggingFace under the Apache 2.0 license. Additionally, submodular optimization was performed using SUBMODLIB, which is an openly accessible library on GitHub at https://github.com/decile-team/submodlib under the MIT license.
@@ Appendix C Task Scaling Curves: Mistral, Falcon
$N^{\prime}$ Data Mix. Mistral Falcon
$25000$ EPM(Baseline-1) 52.69 23.09
EM(Baseline-2) 51.65 21.1
SMART 53.33 22.73
$50000$ EPM(Baseline-1) 53.75 19.86
EM(Baseline-2) 53.16 22.96
SMART 53.8 26.31
$100000$ EPM(Baseline-1) 54.24 20.88
EM(Baseline-2) 54.57 24.85
SMART 53.75 27.22
$200000$ EPM(Baseline-1) 55.25 24.7
EM(Baseline-2) 54.71 24.39
SMART 55.4 30.96
Table 4: Comparison of SMART with baselines for Mistral and Falcon.